If a couple plans to have 5 children, what is the probability that there will be at least one boy? Assume boys and girls are Please answer both questions
Answer #1

From the given information,

Here,

Sample size (n)=5

Probability of success (p)=0.5

Hence,

Using binomial probability calculater,

Binomial Distribution X ~ Bin(n,p) p= 0.5 P(X 2 x) = 0.96875 P(X=x) 0 1 2 3 4 u= E(X) = 2.5 0 = SD(X) = 1.118 m2 = Var(X) = 1

Hence,

Required probability is,

P(X≥1)=0.9688

As this probability very close to 1,

Hence, probability is high enough for the couple to be very confident that they will get at least one boy in 5 children.

Thank you.